OpenStudy (pokemon23):

If 20 is added to a number the result is the same as 55 decreased by four times the number.Find the number. Help me If 20 is added to a number the result is the same as 55 decreased by four times the number.Find the number. Help me @Mathematics

OpenStudy (pokemon23):

20+x=55-4x

OpenStudy (heisenberg):

setup an equation with simple algebra. The "unknown" = the "number."

OpenStudy (pokemon23):

is this correct?

OpenStudy (heisenberg):

Yep that's right.

OpenStudy (pokemon23):

thanks

OpenStudy (heisenberg):

can you solve those together? there are a few techniques for this.

OpenStudy (pokemon23):

let me see

OpenStudy (pokemon23):

20=55-3x

OpenStudy (pokemon23):

75=-3x

OpenStudy (heisenberg):

not quite. to get rid of the x on the left, subtract x to both sides. -4x - (1)x = -5x

OpenStudy (heisenberg):

same with the 55.

OpenStudy (pokemon23):

75=-5x

OpenStudy (heisenberg):

what about: 25 - 55 = ?

OpenStudy (pokemon23):

isn't 20

OpenStudy (pokemon23):

subtract 55

OpenStudy (heisenberg):

it's a negative number.

OpenStudy (pokemon23):

?????????

OpenStudy (heisenberg):

20 - 55 = -35

OpenStudy (heisenberg):

you have to subtract 55 from both sides.

OpenStudy (pokemon23):

where did you get -35

OpenStudy (heisenberg):

20 - 55 = 20 + -55 = -55 + 20 = -35

OpenStudy (pokemon23):

-35=-5x

OpenStudy (heisenberg):

that's right.

OpenStudy (pokemon23):

x=7

OpenStudy (pokemon23):

ty sir?

OpenStudy (heisenberg):

booyah.

OpenStudy (pokemon23):

lol